I Need Help With This Question:
Given the quadratic equation y = 2xSquared + 4x + 5 identify the value of y when x = 0

Answer by jim_thompson5910(35256)

Simply plug in x = 0 (ie replace 'x' with zero), evaluate and simplify.

y = 2x^2 + 4x + 5

y = 2(0)^2 + 4(0) + 5

y = 2(0) + 4(0) + 5

y = 0 + 0 + 5

y = 5

So the value of y is y = 5 when x = 0.